[Brahms-l] Shift report 20040127 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Tue Jan 27 2004 - 07:44:14 EST
------------------------------------------------------------
Date: Tue Jan 27 07:44:14 2004, Shift: 00:00-08:00
Supervisor: Ionut ARSENE, with: 
Log book#15 pages 43 - 43

Shift Summary:
------------------------------------------------------------
1. Beam for 5 hours.

2. Achievements:
 Data
 FS@8deg with D1@A861 and MRS@40deg with D5@A2255
  runs: 9903, 9904, 9905 (570k trig6 and 480k trig3);
 This setting is DONE.
 FS@8deg with D1@A1219 and MRS@40deg with D5@A2255
  runs: 9906, 9907 (~400k trig6 and ~300k trig3).
 
 Pedestals
  run 9901 - Non-TPC pedestals
  run 9902 - TPC pedestals

3. No detector problems. 
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Jan 27 07:44:37 2004

This archive was generated by hypermail 2.1.8 : Tue Jan 27 2004 - 07:44:52 EST